12 men produce 700 watches in 9 working days. How long will it take 18 men to produce the 700 watches
Answer by jorel1380(3719)

If you have 12 men, and 9 working days, then you have 12*9, or 108 man-days.
700/108=6.481 watches per man per day.
700 watches/18 men=38.89 watches per man
38.89/6.481=6 days for the men to produce the watches.
Or:
Since you have 18/12 of the original men, or 3/2; then the watches will take 2/3rds of the original time. 2/3 x 9=6. ☺☺☺☺